 Company just reported Earning Per Share of $50.00 (EPS0=$50)  Management plans to continue to payout 20% of net income in the form of dividends  Before implementing any cost cutting initiatives, management expects earnings will continue to grow at a long term sustainable growth rate of10.0% every year in perpetuity  The company’s cost of equity is 25%

Assume Management does not attempt to cut costs and use the 10% earnings growth rate. Using the Dividend Discount Model for constant growth, the present value of Company’s future growth opportunities (i.e. PVGO) is closest to:

a) $25

b) $27

c) $29

d) $31

e) $33
Answer #1

PVGO = Value of stock – (earnings / cost of equity)

Value of stock = D1 / (Ke - g)

D1 = Do * (1 + g)

D0 = EPS0 * Payout %

D0 = $50 * 20%

D0 = $10

D1 = $10 * (1 + 10%) = $11

Ke = 25%

g = 10%

Value of stock = $11 / (25% - 10%) = $73.3

PVGO = $73.3 - ($11 / 25%)

PVGO = $73.3 - $44

PVGO = $29.3

Correct choice - c - $29